Muscarella Inc. has the following balance sheet and income statement data:
Cash $ 14,000 Accounts payable $ 42,000
Receivables 70,000 Other current liabilities 28,000
Inventories 210,000 Total CL $ 70,000
Total CA $294,000 Long-term debt 70,000
Net fixed assets 126,000 Common equity 280,000
Total assets $420,000 Total liab. and equity $420,000
Sales $280,000
Net income $ 21,000
The new CFO thinks that inventories are excessive and could be lowered sufficiently to cause the current ratio to equal the industry average, 2.70, without affecting either sales or net income. Assuming that inventories are sold off and not replaced to get the current ratio to the target level, and that the funds generated are used to buy back common stock at book value, by how much would the ROE change?
1
Expert's answer
2012-07-19T07:28:44-0400
We can use the next formula:
Current ratio= Current assets/ Current Liabilities
ROE = Net Income / Shareholder’s Equity
ΔROE = ROE2 - ROE1
So
ROE = 21000/280 000 = 0,075 or 7,5%
Inv at target CR = 105 000 $
We have decreasing in common equity and inventory, it is 105 000$
So the new common equity will be 175 000$
ROE2 = 0,12 or 12%
ΔROE = 12% - 7,5% = 4,5%